OpenStudy (lilai3):

At the craft store, a paint brush costs $0.79, and a small bottle of paint costs $0.89. What combination of paint brushes and bottles of paint could you buy for exactly $4.15? (No tax included. NOT A TAX PROBLEM.)

OpenStudy (anonymous):

this looks like a guess and check problem to me

OpenStudy (anonymous):

\(1\times .79+4\times .89=4.35\) nope lets try again

OpenStudy (anonymous):

\(2\times .79+3\times .89=4.25\) ok not that one either

OpenStudy (anonymous):

bet we get it on the next try though...

OpenStudy (anonymous):

oh yeah, next one works

OpenStudy (lilai3):

hello any1? i waited for like one hour and a half already -_-

OpenStudy (anonymous):

did you not understand my answer?

OpenStudy (anonymous):

maybe it was not clear. method is to guess and check to see if it works started with one paint brush and four bottles , got a total cost of \[1\times .79+4\times .89=4.35\] which is too big, then tried 2 paint brushes and 3 bottle for a total cost of \[2\times .79+3\times .89=4.25\] again too big so what would the next one in this list be?

OpenStudy (lilai3):

i actually don't know because there are unlimited answers

OpenStudy (anonymous):

did you try the next one?

OpenStudy (lilai3):

i didn't because my point is, there are unlimited numbers to plug into the equation

OpenStudy (anonymous):

that is i suppose technically true, that is you could try any number of brushes and paint, but not any will add to $4.15

OpenStudy (anonymous):

for example one million of each would be silly

OpenStudy (lilai3):

yes that would be my point, but it would be unlimited money LOL :P yes that's my point, but what i'm trying to say is, you can do 5 or 6 or 7 or 8 or 9 or 10 or 11 or 12 or 13 or 14 or 15 and so on and so forth

OpenStudy (anonymous):

i understand that you have only one equation and two unknowns, and so if the number of brushes and the number of cans could be any number at all, then you would have an infinite number of solutions to \(.79x+.89y=4.15\)

OpenStudy (anonymous):

but in this case \(x\) and \(y) must be whole numbers, not any number

OpenStudy (shane_b):

There isn't an unlimited number of answers because there's only a few combinations that won't exceed $4.15 and x and y must be whole numbers.

OpenStudy (anonymous):

and of course in your examples above, none of the work because they add up to more than $4.15

OpenStudy (anonymous):

an equation with only integer solutions goes by the fancy name of "diophantine" equations you are asked here for whole number solutions, and the way to find them is, well, to find them guess and check, see what works

OpenStudy (lilai3):

so do i use the equation 0.79x +0.89y=4.15?